# Retrieval-Enhanced Knowledge Editing for Multi-Hop Question Answering in Language Models

### v1

@@ 3.1. Retrieval-Augmented Editing
A naïve approach is using similarity-based search to retrieve edited facts similar to target question $q$ (Wang et al., 2023a; Han et al., 2023; Zhong et al., 2023). These facts are then integrated into a prompt template for editing via in-context learning: $f_{\theta}^{\prime}(q)=f_{\theta}(T_{e}(q,\{\delta_{1}^{\prime},\delta_{2}^{\prime},...\delta_{K}^{\prime}\})),$ where $T_{e}$ is the editing template. For example, $T_{e}(\cdot)$ can be made as ”Given fact: $\{\delta^{\prime}\}$, $\{q\}$”. The Top-$K$ nearest edited facts to question $q$ in the embedding space are denoted as $\{\delta_{1}^{\prime},\delta_{2}^{\prime},...\delta_{K}^{\prime}\}={\text{Top-}K}_{\delta\in\Delta}\,\text{sim}(g_{z}(\delta),g_{z}(q))$, where $\text{sim}(\cdot)$ denotes the similarity function and $g_{z}$ is an embedding model. However, edited facts $\Delta_{q}$ needed to answer $q$ are hard to retrieve by this approach since they usually contain entities different from $q$, which will result in a low similarity score in a large bank $\Delta$ (e.g., in Table 1, United States in $\delta_{3}^{\prime}$, but not in $q$).
To address this problem, we propose edited fact chain extraction to obtain $G^{*}_{q}$. Inherently, each $G^{*}_{q}$ forms a connected knowledge graph (KG) (Zhong et al., 2023). Such KGs can be retrieved by iteratively traversing links from one entity to another. Take $G^{*}_{q}=\{\delta^{\prime}_{1},\delta_{2},\delta^{\prime}_{3}\}$ in Table 1 as an example. It is composed of two edited facts $\delta^{\prime}_{1},\delta^{\prime}_{3}$ and one unedited fact $\delta_{2}$. We can observe that the question entity: (Harry Potter) is the head entity $h_{1}$ in $\delta_{1}^{\prime}=\{h_{1},t_{1},t_{1}^{\prime}\}$, and the edited tail entity $t^{\prime}_{1}$: (Stephen King) is also the head entity $h_{2}$ in the next fact $\delta_{2}=\{h_{2},r_{2},t_{2}\}$. Moreover, for each subsequent fact in the chain, its head entity is always the tail entity of the previous fact. By effectively retrieving the KG that represents the fact chain $G^{*}_{q}$, we are able to capture all the edited factual triplets $\Delta_{q}=\{\delta_{i}^{\prime},\delta_{j}^{\prime},...\delta_{k}^{\prime}\}$. In light of this, we define our retrieval-augmented editing as:
@@ 3.2.1. External Knowledge Graph for Subgraph Retrieval
According to our previous discussion, we aim to retrieve the fact chain $G^{*}_{q}$ for model editing. However, it is worth noting that $G^{*}_{q}$ consists of both edited and unedited facts, whereas the unedited facts do not exist in our edited fact bank $\Delta$ by default. To effectively incorporate both types of facts into our retrieval process, we propose integrating all edited facts into an external knowledge graph $\mathcal{G}$. By selecting a comprehensive KG such as WikiData (Vrandečić and Krötzsch, 2014), the new graph $\mathcal{G}^{*}$ will encompass both unedited and edited facts. It complements our edited fact bank $\Delta$ and connects different entities. Besides, the external knowledge graph provides extra factual knowledge that can enhance language to output correct answers.
Specifically, given the edits $\Delta=\{\delta_{1}^{\prime},...,\delta_{n}^{\prime}\}$ and an external $\mathcal{G}$, we consider two types of operations to combine them. (1) Modifying existing facts: If the original fact appears in $\mathcal{G}$, i.e., $(h,r,t)\in\mathcal{G}$, we will modify the KG according to the edits, so $\mathcal{G}^{*}=(h,r,t^{\prime})\cup\mathcal{G}\setminus(h,r,t)$. (2) Adding new facts: If the original fact does not appear in $\mathcal{G}$, i.e., $(h,r,t)\notin\mathcal{G}$, then we append the modified fact to the KG, so $\mathcal{G}^{*}=(h,r,t^{\prime})\cup\mathcal{G}$. Next, given a question $q$, we retrieve a subgraph $G_{S}$ from $\mathcal{G}^{*}$, so that $G_{S}\subset\mathcal{G}^{*}$. Our goal is to ensure that $G_{S}$ contains fact chains corresponding to $q$, i.e., $G^{*}_{q}\subseteq G_{S}$.
@@ 3.2.2. Mutual Information based Retrieval Objective
(2) $\max_{G_{S}}I\left(Q,G_{S}\right)=H(Q)-H\left(Q\mid G=G_{S}\right).$
Given a fixed question set $Q$, its Shannon entropy $H(Q)$ remains constant. Therefore, maximizing the mutual information $I\left(Q,G_{S}\right)$ is equivalent to minimizing the conditional entropy $H\left(Q\mid G=G_{S}\right)$. Thus, we focus on optimizing the following objective:
(3) $\max_{G_{S}}I\left(Q,G_{S}\right)=\min_{G_{S}}H\left(Q\mid G=G_{S}\right)$
@@
In practice, quantifying $p(q|G=G_{S})$ poses a significant challenge due to its computational complexity. This complexity arises as there are numerous subgraph candidates $G_{S}$ within the entire knowledge graph, making it prohibitively expensive to exhaustively search for the optimal subgraph. To circumvent this issue, we first replace the intractable term $p(q|G=G_{S})$ with $\frac{p(q,G=G_{S})}{p(G=G_{S})}$. Then, suppose we consider one question each time, where $Q={q}$, we can reformulate the objective as:
@@ 3.2.3. Probabilities Estimation
This decomposition allows us to initially focus on estimating the $p(r_{1}|q,h_{1})$. Specifically, the head entity $h_{1}$ is determined if $q$ is given, since we assume $h_{1}$ is mentioned in question $q$. Candidate relations for $r_{1}$ can also be selected from the edited KG. Practically, we can estimate the probability $p(r_{1}|q,h_{1})$ for each candidate relation using an auto-regressive language model $f_{\phi}$ (Radford et al., 2019; Wu et al., 2023):
@@
where $f_{\phi}$ is the predicted word probability, and $w_{q},w_{h_{1}},w_{r_{1}}$ denote words in the question $q$, head entity $h_{1}$, and relation $r_{1}$, respectively. We can employ open-source LLMs like GPT-2 (Radford et al., 2019) for this estimation. Please note that, the model $f_{\theta}$ being edited does not need to be the same model used for probability estimation, making our method applicable even for editing proprietary LLMs. With a specific input context $\{q,h_{1}\}$, the language model will assign different probabilities to each relation based on its contextual understanding and reasoning ability.
@@
Then, for the last term $p(q,h_{1})/p(h_{1})$ in Equation (6), based on Bayes’ theorem, we can transform it into $p(q,h_{1})/p(h_{1})=p(q|h_{1})$, which is a constant value given a specific question $q$. We can also apply model $f_{\phi}$ to estimate this conditional probability. Now, since we are able to estimate every term in Equation (6) and (5), we can effectively identify the subgraph that yields the maximum Mutual Information. Additionally, we utilize the beam search technique (SCIENCE, 1977) to expedite the computational process, eliminating the necessity for exhaustively traversing all connected nodes. Further details on our implementations are provided in the Appendix B. In this work, we treat $n$ as a hyperparameter since the number of hops required to answer a question is unknown in advance. To ensure thorough exploration, $n$ is assigned a large value, enabling an extensive search range. However, this approach will also introduce irrelevant information in the retrieved subgraph $G_{S}$, which can potentially mislead the language model to generate undesired answers (Li et al., 2024, 2023). Thus, in the next section, we will discuss how to mitigate this problem.
@@ 3.3.1. Editing Uncertainty
(11) $H(Y|X=x)=-\sum_{y}p(y|x)\log_{2}(p(y|x)),$
where $y$ represents each possible answer generated by the language model, and $x=\{q,G_{S}\}$ is the model input composed of the question $q$ and facts $G_{S}$. A higher entropy value $H(Y|X=x)$ signifies less confidence in the answer, reflecting greater uncertainty. Conversely, a lower entropy value indicates higher confidence and less uncertainty. Ideally, if input facts $G_{S}$ are exactly the edited question fact chain $G^{*}_{q}$, i.e., $G_{S}=G^{*}_{q}$, then the model output should exhibit maximum confidence with minimal entropy, since $G^{*}_{q}$ contains the precise knowledge to answer question $q$. In the next part, we conduct empirical experiments to verify this assumption.
In our experiments, GPT-J (6B) (Wang and Komatsuzaki, 2021) is chosen as the base language model. We select 1000 instances for each of 2, 3, and 4-hop questions from the MQUAKE-CF dataset (Zhong et al., 2023) for testing. The MQUAKE-CF dataset comprises multi-hop questions that are based on real-world facts, where the edited facts are counterfactual, meaning they do not exist in actual real-world scenarios. An example of such a question with an edit is provided in Table 1.
Our experiment seeks to identify the fact set $G^{\prime}_{S}$ that, when used as model input, yields the lowest output entropy, indicating minimal editing uncertainty. Our first step is to construct different fact set candidates. We begin with the first fact $\delta_{1}$ in the fact chain $G^{*}_{q}$ as our initial fact set $G^{\prime}_{S}$. Then, we add each subsequent fact from the chain until the $G^{\prime}_{S}$ encompasses the entire fact chain $G^{*}_{q}$. After that, we introduce unrelated facts $\hat{\delta}$ into the set. This process is repeated until $G^{\prime}_{S}$ contains six elements. Finally, we build a prefix set $\bar{G}_{q}$ with all the six subsets $G^{\prime}_{S}$. Specifically, for a $4$-hop question, we have $\bar{G}_{q}=\{\{\delta_{1}\},\{\delta_{1},\delta_{2}\},\{\delta_{1},\delta_{2},\delta_{3}\},...,\{\delta_{1},\delta_{2},\delta_{3},\delta_{4},\hat{\delta}_{5},\hat{\delta}_{6}\}\}$, where $\delta_{1},\delta_{2},\delta_{3},\delta_{4}\in G^{*}_{q}$ and $\hat{\delta}_{5},\hat{\delta}_{6}$ are two irrelevant facts. Finally, we conduct in-context editing using each subset $G^{\prime}_{S}$ from the prefix set $\bar{G}_{q}$ with editing template $T_{e}$:”Given fact: $\{G^{\prime}_{S}\}$, $\{q\}$?”. In our experiments, for computational simplicity, we consider model output $y$ to be each of the next predicted word. We report the entropy over all the words in the vocabulary as the editing uncertainty.
@@
(a) Fact chain $G^{*}_{q}$ with redundant knowledge.
(b) Random fact chain without useful knowledge.
Figure 3. Distribution of normalized model editing entropy with different fact subsets as input. A lower normalized entropy on a given subset indicates that the model is more confident in answering the question with the given facts. Subset 1 includes the first fact $\{\delta_{1}\}$, Subset 2 includes the first two facts $\{\delta_{1},\delta_{2}\}$, and so on. Figure 3(a) shows that the entropy is significantly lower if the subset contains exactly the entire fact chain of the question (e.g., Subset 2 for 2-hop questions).
@@ 3.4. Theoretical Justification
In this subsection, we theoretically justify that the facts collected by our retrieval objective (Eq. (2)) are effective in performing model editing with in-context learning. To begin with, we discuss what kinds of input can effectively activate in-context learning. Then, we explore how to build such effective input for model editing.
@@ 4. Experiments
We conduct experiments to answer the following questions. Q1: Does RAE successfully edit model output? Q2: How does our retrieval strategy perform compared with other retrieval methods? Q3: Does our proposed pruning technique remove redundant facts from the retrieved facts? Q4: Does RAE work for propriety LLMs?
@@ 4.1.1. Language Models
We evaluate RAE across various kinds of language models in different sizes and families, including GPT-2 (1.3B)  (Radford et al., 2019), GPT-J (6B) (Wang and Komatsuzaki, 2021), Falcon (7B) (Almazrouei et al., 2023), Vicuna (7B) (Chiang et al., 2023), and Llama2-chat (7B) (Touvron et al., 2023a). Among them, GPT-2, GPT-J, and Falcon are pre-trained language models without instruction tuning (Ouyang et al., 2022; Chung et al., 2022), while Vicuna is an instruction-tuned variation of Llama1 (Touvron et al., 2023b) and Llama2-chat is the instruction-tuned version of Llama2. Instruction-tuned models (Vicuna and Llama2-chat) are expected to better follow the instructions in the prompt compared to native pre-trained models (GPT-2, GPT-J, and Falcon). We include both kinds of models to verify the effectiveness of the proposed methods.

### v2

@@ 3.1. Retrieval-Augmented Editing
A naïve edit approach uses similarity-based search to retrieve edited facts similar to target question $q$ (Wang et al., 2023; Han et al., 2023; Zhong et al., 2023). These facts are then integrated into a prompt template for editing via in-context learning: $f_{\theta}^{\prime}(q)=f_{\theta}(T_{e}(q,\{\delta_{1}^{\prime},\delta_{2}^{\prime},...\delta_{K}^{\prime}\})),$ where $T_{e}$ is the editing template. For example, $T_{e}(\cdot)$ can be made as ”Given fact: $\{\delta^{\prime}\}$, $\{q\}$”. The Top-$K$ nearest edited facts to question $q$ in the embedding space are denoted as $\{\delta_{1}^{\prime},\delta_{2}^{\prime},...\delta_{K}^{\prime}\}={\text{Top-}K}_{\delta\in\Delta}\,\text{sim}(g_{z}(\delta),g_{z}(q))$, where $\text{sim}(\cdot)$ denotes the similarity function and $g_{z}$ is an embedding model. However, the edited facts $\Delta_{q}$ needed to answer $q$ are difficult to retrieve by this approach since they usually contain entities different from $q$, which will result in a low similarity score in a large bank $\Delta$ (e.g., in Table 1, United Kingdom in $\delta_{3}^{\prime}$, but not in $q$).
To address this problem, we propose edited fact chain extraction to obtain $G^{*}_{q}$. Inherently, each $G^{*}_{q}$ forms a connected knowledge graph (KG) (Zhong et al., 2023). Such KGs can be retrieved by iteratively traversing links from one entity to another. Take $G^{*}_{q}=\{\delta^{\prime}_{1},\delta_{2},\delta^{\prime}_{3}\}$ in Table 1 as an example. It is composed of two edited facts $\delta^{\prime}_{1},\delta^{\prime}_{3}$ and one unedited fact $\delta_{2}$. We can observe that the question entity: (Misery) is the head entity $h_{1}$ in $\delta_{1}^{\prime}=\{h_{1},t_{1},t_{1}^{\prime}\}$, and the edited tail entity $t^{\prime}_{1}$: (Richard Dawkins) is also the head entity $h_{2}$ in the next fact $\delta_{2}=\{h_{2},r_{2},t_{2}\}$. Moreover, for each subsequent fact in the chain, its head entity is always the tail entity of the previous fact. By effectively retrieving the KG that represents the fact chain $G^{*}_{q}$, we are able to capture all the edited factual triplets $\Delta_{q}=\{\delta_{i}^{\prime},\delta_{j}^{\prime},...\delta_{k}^{\prime}\}$. In light of this, we define our retrieval-augmented editing as:
@@ 3.2.1. External Knowledge Graph for Subgraph Retrieval
According to our previous discussion, we aim to retrieve the fact chain $G^{*}_{q}$ for model editing. Additionally, it is worth noting that $G^{*}_{q}$ consists of both edited and unedited facts. However, the unedited facts are not included in our edited fact bank $\Delta$ by default. To effectively incorporate both types of facts into our retrieval process, we propose integrating all edited facts into an external knowledge graph $\mathcal{G}$. By selecting a comprehensive KG such as WikiData (Vrandečić and Krötzsch, 2014), the new graph $\mathcal{G}^{*}$ will encompass both unedited and edited facts. It complements our edited fact bank $\Delta$ and connects different entities. Besides, the external knowledge graph provides extra factual knowledge that can enhance language to output correct answers.
Specifically, given the edits $\Delta=\{\delta_{1}^{\prime},...,\delta_{n}^{\prime}\}$ and an external $\mathcal{G}$, we consider two types of operations to combine them. (1) Modifying existing facts: If the original fact appears in $\mathcal{G}$, i.e., $(h,r,t)\in\mathcal{G}$, we will modify the KG according to the edits, so $\mathcal{G}^{*}=(h,r,t^{\prime})\cup\mathcal{G}\setminus(h,r,t)$. (2) Adding new facts: If the original fact does not appear in $\mathcal{G}$, i.e., $(h,r,t)\notin\mathcal{G}$, then we append the modified fact to the KG, so $\mathcal{G}^{*}=(h,r,t^{\prime})\cup\mathcal{G}$. Next, given a question $q$, we retrieve a subgraph $G_{S}$ from $\mathcal{G}^{*}$, so that $G_{S}\subset\mathcal{G}^{*}$. Our goal is to ensure that $G_{S}$ contains fact chains of $q$, i.e., $G^{*}_{q}\subseteq G_{S}$.
@@ 3.2.2. Mutual Information based Retrieval Objective
(2) $\max_{G_{S}}I\left(Q;G_{S}\right)=H(Q)-H\left(Q\mid G=G_{S}\right).$
Given a fixed question set $Q$, its Shannon entropy $H(Q)$ is constant. Therefore, maximizing the mutual information $I\left(Q;G_{S}\right)$ is equivalent to minimizing the conditional entropy $H\left(Q\mid G=G_{S}\right)$. Thus, we optimize the following objective:
(3) $\max_{G_{S}}I\left(Q;G_{S}\right)=\min_{G_{S}}H\left(Q\mid G=G_{S}\right)$
@@
In practice, quantifying $p(q|G=G_{S})$ is challenging due to its computational complexity. This is because there are numerous subgraph candidates $G_{S}$ within the entire knowledge graph, making it prohibitively expensive to exhaustively search for the optimal one. To circumvent this issue, we first replace the intractable term $p(q|G=G_{S})$ with $\frac{p(q,G=G_{S})}{p(G=G_{S})}$. Then, suppose we consider one question each time, where $Q={q}$, the objective is reformulated as:
@@ 3.2.3. Probabilities Estimation
This decomposition allows us to initially focus on estimating the $p(r_{1}|q,h_{1})$. Specifically, the head entity $h_{1}$ is determined if $q$ is given, since we assume $h_{1}$ is mentioned in question $q$. Candidate relations for $r_{1}$ can also be selected from the edited KG. Practically, we can estimate the probability $p(r_{1}|q,h_{1})$ for each candidate relation using an auto-regressive language model $f_{\phi}$ (Radford et al., 2019; Wu et al., 2024):
@@
where $f_{\phi}$ is the predicted word probability, and $w_{q},w_{h_{1}},w_{r_{1}}$ denote the words in question $q$, head entity $h_{1}$, and relation $r_{1}$, respectively. We can employ open-source LLMs like GPT-2 (Radford et al., 2019) for this estimation. Please note that, the model $f_{\theta}$ being edited does not need to be the same model used for probability estimation, making our method applicable even for editing proprietary LLMs. With a specific input context $\{q,h_{1}\}$, the language model will assign different probabilities to each relation based on its contextual understanding and reasoning ability.
@@
Then, for the last term $p(q,h_{1})/p(h_{1})$ in Equation (6), based on Bayes’ theorem, we can transform it into $p(q,h_{1})/p(h_{1})=p(q|h_{1})$, which is a constant value given a specific question $q$. We can also apply model $f_{\phi}$ to estimate this conditional probability. Now, since we are able to estimate every term in Equation (5) and (6), we can effectively identify the subgraph that yields the maximum Mutual Information. Additionally, we utilize beam search (SCIENCE, 1977) to expedite the computational process, eliminating the necessity for exhaustively traversing all connected nodes. Further details on our implementations are provided in the Appendix B. In this work, we treat $n$ as a hyperparameter since the number of hops required to answer a question is unknown in advance. To ensure thorough exploration, $n$ is assigned a large value. However, this approach will also introduce irrelevant information in the retrieved subgraph $G_{S}$, which can potentially mislead the language model to hallucinate and generate undesired answers (Li et al., 2024, 2023). In the next section, we will discuss how to mitigate this problem.
(a) Fact chain $G^{*}_{q}$ with redundant knowledge.
(b) Random fact chain without useful knowledge.
Figure 3. Distribution of normalized model editing entropy with different fact subsets as input. A lower normalized entropy indicates that the model is more confident in answering the question with the given facts. ”Subset 1” includes the first fact $\{\delta_{1}\}$, ”Subset 2” includes the first two facts $\{\delta_{1},\delta_{2}\}$, and so on. Figure 3(a) shows that the entropy is significantly lower if the subset contains exactly the entire fact chain of the question (e.g., Subset 2 has low entropy for 2-hop questions).
@@ 3.3.1. Editing Uncertainty
(11) $H(Y|X=x)=-\sum_{y}p(y|x)\log_{2}p(y|x),$
where $y$ represents each possible answer generated by the language model, and $x=\{q,G_{S}\}$ is the model input composed of the question $q$ and facts $G_{S}$. A higher entropy value $H(Y|X=x)$ means less confidence in the answer, reflecting greater uncertainty. In contrast, a lower entropy value indicates higher confidence and less uncertainty. Ideally, if input facts $G_{S}$ are exactly the edited question fact chain $G^{*}_{q}$, i.e., $G_{S}=G^{*}_{q}$, then the model output should exhibit maximum confidence with minimal entropy, since $G^{*}_{q}$ contains the precise knowledge to answer question $q$. In the next part, we conduct empirical experiments to verify this assumption.
In our experiments, we choose GPT-J (6B) (Wang and Komatsuzaki, 2021) as the base language model. We select 1000 instances for each of the 2, 3 and 4-hop questions from the MQUAKE-CF dataset (Zhong et al., 2023) for testing. The MQUAKE-CF dataset comprises multi-hop questions that are based on real-world facts, where the edited facts are counterfactual, meaning they do not exist in actual real-world scenarios. An example of such a question with an edit is provided in Table 1.
Our experiment seeks to identify the fact set $G^{\prime}_{S}$ that, when used as model input, yields the lowest output entropy (i.e., minimal editing uncertainty). Our first step is to construct different fact set candidates. We begin with the first fact $\delta_{1}$ in the fact chain $G^{*}_{q}$ as our initial fact set $G^{\prime}_{S}$. Then, we add each subsequent fact from the chain until the $G^{\prime}_{S}$ encompasses the entire fact chain $G^{*}_{q}$. After that, we insert unrelated facts $\hat{\delta}$ into the set. In this experiment, the process is repeated until $G^{\prime}_{S}$ contains six elements, where we build a prefix set $\bar{G}_{q}$ with all the six subsets $G^{\prime}_{S}$. For example, for a $4$-hop question, we have $\bar{G}_{q}=\{\{\delta_{1}\},\{\delta_{1},\delta_{2}\},\{\delta_{1},\delta_{2},\delta_{3}\},...,\{\delta_{1},\delta_{2},\delta_{3},\delta_{4},\hat{\delta}_{5},\hat{\delta}_{6}\}\}$, where $\delta_{1},\delta_{2},\delta_{3},\delta_{4}\in G^{*}_{q}$ and $\hat{\delta}_{5},\hat{\delta}_{6}$ are two irrelevant facts. Finally, we conduct in-context editing using each subset $G^{\prime}_{S}$ from $\bar{G}_{q}$ with editing template $T_{e}$:”Given fact: $\{G^{\prime}_{S}\}$, $\{q\}$?”. In our experiment, we consider model output $y$ to be each of the next predicted word. We report the entropy over all the words in the vocabulary as the editing uncertainty.
@@ 3.4. Theoretical Justification
In this subsection, we theoretically justify that the facts collected by our retrieval objective Eq. (2) are effective in performing model editing with in-context learning. To begin with, we discuss what kinds of input can effectively activate in-context learning. Then, we explore how to build such effective input for model editing.
@@ 4. Experiments
We conduct experiments to answer the following questions. Q1: How effective is RAE in editing LLM output? Q2: How does our retrieval strategy perform compared to other retrieval methods? Q3: Does our proposed pruning technique remove redundant facts from the retrieved facts? Q4: Does RAE work for propriety LLMs?
@@ 4.1.1. Language Models
We evaluate RAE across various kinds of language models in different sizes and families, including GPT-2 (1.5B)  (Radford et al., 2019), GPT-J (6B) (Wang and Komatsuzaki, 2021), Falcon (7B) (Almazrouei et al., 2023), Vicuna (7B) (Chiang et al., 2023), and Llama2-chat (7B) (Touvron et al., 2023a). Among them, GPT-2, GPT-J, and Falcon are pre-trained language models without instruction tuning (Ouyang et al., 2022; Chung et al., 2022), while Vicuna is an instruction-tuned variation of Llama1 (Touvron et al., 2023b) and Llama2-chat is the instruction-tuned version of Llama2. Instruction-tuned models (Vicuna and Llama2-chat) are expected to better follow the instructions in the prompt compared to native pre-trained models (GPT-2, GPT-J, and Falcon). We include both kinds of models to verify the effectiveness of the proposed methods.